# MCGS-SLAM: A Multi-Camera SLAM Framework Using Gaussian Splatting for High-Fidelity Mapping

### v1

@@ IV-B Rendering Results Study
Method Metric 100613 106762 134763 152706 NICER- SLAM [14] PSNR $\uparrow$ 12.91 13.09 8.79 11.32 SSIM  $\uparrow$ 0.498 0.587 0.330 0.611 LPIPS $\downarrow$ 0.695 0.626 0.791 0.754 GLORIE- SLAM [15] PSNR $\uparrow$ 25.78 27.35 25.71 24.90 SSIM  $\uparrow$ 0.916 0.918 0.883 0.878 LPIPS $\downarrow$ 0.282 0.272 0.365 0.338 MonoGS [17] PSNR $\uparrow$ 20.58 22.31 21.41 22.34 SSIM  $\uparrow$ 0.674 0.741 0.620 0.784 LPIPS $\downarrow$ 0.607 0.503 0.625 0.641 DROID- Splat [27] PSNR $\uparrow$ 26.77 27.21 26.20 25.92 SSIM  $\uparrow$ 0.829 0.864 0.792 0.782 LPIPS $\downarrow$ 0.273 0.281 0.376 0.512 Photo- SLAM [28] PSNR $\uparrow$ 19.03 20.26 21.28 20.84 SSIM  $\uparrow$ 0.640 0.712 0.624 0.759 LPIPS $\downarrow$ 0.527 0.440 0.471 0.466 MCGS- SLAM (Ours) PSNR $\uparrow$ 27.09 27.70 27.20 28.45 SSIM  $\uparrow$ 0.830 0.819 0.813 0.797 LPIPS $\downarrow$ 0.223 0.262 0.233 0.330
TABLE I: Appearance reconstruction comparison of different methods on four scenes of the Waymo dataset [22] (Real-World Dataset). MCGS-SLAM achieves the best PSNR and LPIPS results, highlighted as first, second and third.
@@
Figure 4: MCGS-SLAM produces faithful and complete reconstructions on AirSim [26] (Synthetic Dataset), demonstrated in the Factory (a) and Village (b) environments, where it preserves structural details and textures that monocular methods often fail to capture.
@@
Method   Metric   Garden   Factory   Village NICER- SLAM [14] PSNR $\uparrow$ 12.30 9.84 11.18 SSIM  $\uparrow$ 0.450 0.332 0.504 LPIPS $\downarrow$ 0.801 0.690 0.653 GLORIE- SLAM [15] PSNR $\uparrow$ 24.50 23.39 17.56 SSIM  $\uparrow$ 0.849 0.888 0.494 LPIPS $\downarrow$ 0.351 0.346 0.712 MonoGS [17] PSNR $\uparrow$ 25.59 21.45 21.39 SSIM  $\uparrow$ 0.766 0.760 0.689 LPIPS $\downarrow$ 0.258 0.175 0.444 DROID- Splat [27] PSNR $\uparrow$ 24.12 26.50 17.25 SSIM  $\uparrow$ 0.822 0.898 0.669 LPIPS $\downarrow$ 0.242 0.107 0.652 Photo- SLAM [28] PSNR $\uparrow$ 25.47 28.38 26.77 SSIM  $\uparrow$ 0.775 0.923 0.805 LPIPS $\downarrow$ 0.156 0.041 0.205 MCGS- SLAM (Ours) PSNR $\uparrow$ 29.36 28.37 28.10 SSIM  $\uparrow$ 0.879 0.924 0.853 LPIPS $\downarrow$ 0.126 0.083 0.219
@@
Figure 5: Tracking performance on the Oxford Spires Dataset [25], evaluated across four representative sequences. Ground truth trajectories are compared against Splat-SLAM [19], HI-SLAM2 [20], and our MCGS-SLAM. MCGS-SLAM remains closely aligned with ground truth across all sequences, usually achieving the lowest ATE RMSE values and demonstrating the robustness and accuracy of our multi-camera framework in large-scale outdoor environments.

### v2

@@ IV-B Rendering Results Study
Method   Metric   100613   132384   134763   152706   158686   153495   106762   163453   Avg. NICER-SLAM [14] PSNR $\uparrow$ 12.91 15.48 8.79 11.32 11.68 13.25 13.09 16.41 12.87 SSIM  $\uparrow$ 0.498 0.775 0.330 0.611 0.438 0.541 0.587 0.712 0.562 LPIPS $\downarrow$ 0.695 0.518 0.791 0.754 0.686 0.691 0.626 0.657 0.677 GLORIE-SLAM [15] PSNR $\uparrow$ 25.78 25.52 25.71 24.90 25.09 23.79 27.35 23.72 25.23 SSIM  $\uparrow$ 0.916 0.902 0.883 0.878 0.908 0.891 0.918 0.903 0.900 LPIPS $\downarrow$ 0.282 0.287 0.365 0.338 0.291 0.309 0.272 0.279 0.303 MonoGS [17] PSNR $\uparrow$ 20.58 23.53 21.41 22.34 21.87 21.08 22.31 19.41 21.57 SSIM  $\uparrow$ 0.674 0.862 0.620 0.784 0.684 0.772 0.741 0.753 0.737 LPIPS $\downarrow$ 0.607 0.421 0.625 0.641 0.514 0.646 0.503 0.657 0.577 DROID-Splat [27] PSNR $\uparrow$ 26.77 25.02 26.20 25.92 26.81 24.01 27.21 23.02 25.62 SSIM  $\uparrow$ 0.829 0.823 0.792 0.782 0.850 0.748 0.864 0.720 0.801 LPIPS $\downarrow$ 0.273 0.384 0.376 0.512 0.297 0.482 0.281 0.451 0.382 Photo-SLAM [28] PSNR $\uparrow$ 19.03 20.49 21.28 20.84 21.44 18.14 20.26 19.12 20.08 SSIM  $\uparrow$ 0.640 0.824 0.624 0.759 0.674 0.726 0.712 0.758 0.715 LPIPS $\downarrow$ 0.527 0.307 0.471 0.466 0.367 0.453 0.440 0.444 0.434 MCGS-SLAM (Ours) PSNR $\uparrow$ 27.09 26.26 27.20 28.45 21.91 26.48 27.70 26.92 26.50 SSIM  $\uparrow$ 0.830 0.826 0.813 0.797 0.682 0.813 0.819 0.829 0.801 LPIPS $\downarrow$ 0.223 0.284 0.233 0.330 0.547 0.231 0.262 0.234 0.293
TABLE I: Appearance reconstruction comparison of different methods on 8 scenes of the Waymo dataset [22] (Real-World Dataset). MCGS-SLAM achieves the best PSNR and LPIPS results, highlighted as first, second and third.
@@
Figure 4: MCGS-SLAM produces faithful and complete reconstructions on AirSim [26] (Synthetic Dataset).
@@
Method Metric Garden Factory Village   Avg. NICER- SLAM [14] PSNR $\uparrow$ 12.30 9.84 11.18 11.11 SSIM  $\uparrow$ 0.450 0.332 0.504 0.429 LPIPS $\downarrow$ 0.801 0.690 0.653 0.715 GLORIE- SLAM [15] PSNR $\uparrow$ 24.50 23.39 17.56 21.82 SSIM  $\uparrow$ 0.849 0.888 0.494 0.744 LPIPS $\downarrow$ 0.351 0.346 0.712 0.470 MonoGS [17] PSNR $\uparrow$ 25.59 21.45 21.39 22.81 SSIM  $\uparrow$ 0.766 0.760 0.689 0.738 LPIPS $\downarrow$ 0.258 0.175 0.444 0.292 DROID- Splat [27] PSNR $\uparrow$ 24.12 26.50 17.25 22.62 SSIM  $\uparrow$ 0.822 0.898 0.669 0.796 LPIPS $\downarrow$ 0.242 0.107 0.652 0.334 Photo- SLAM [28] PSNR $\uparrow$ 25.47 28.38 26.77 26.87 SSIM  $\uparrow$ 0.775 0.923 0.805 0.834 LPIPS $\downarrow$ 0.156 0.041 0.205 0.134 MCGS- SLAM (Ours) PSNR $\uparrow$ 29.36 28.37 28.10 28.64 SSIM  $\uparrow$ 0.879 0.924 0.853 0.885 LPIPS $\downarrow$ 0.126 0.083 0.219 0.143
@@ IV-C Tracking Results Study
Figure 5: Tracking performance on the Oxford Spires Dataset [25], evaluated across 4 representative sequences. Ground truth trajectories are compared against Splat-SLAM [19], HI-SLAM2 [20], and our MCGS-SLAM. MCGS-SLAM remains closely aligned with ground truth across all sequences, usually achieving the lowest ATE RMSE values and demonstrating the robustness and accuracy of our multi-camera framework in large-scale outdoor environments.